# Upper Approximation Bounds for Neural Oscillators

### v1

@@ 2 Neural Oscillator
Assumption 6: Since ${{\Phi}}=[\Phi_{1},\Phi_{2}....,\Phi_{q}]$ is a causal and uniformly continuous operator mapping from $C_{0}([0,T];{\mathbb{R}}^{p})$ to $C_{0}([0,T];{\mathbb{R}}^{q})$, its $j^{\mathrm{th}}$ element $\Phi_{j}$ has a monotonic continuous modulus of continuity $\phi_{\Phi_{j}}(\cdot):[0,+\infty)\to[0,+\infty)$ with $\phi_{\Phi_{j}}(r)\to 0$ as $r\to 0$, such that ${\left\lVert\Phi_{j}[\mathbf{u}_{1}(\tau)](t)-\Phi_{j}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}$ $\leq\phi_{\Phi_{j}}[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}}]$ for arbitrary $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in C_{0}([0,T];{\mathbb{R}^{p}})$. $\phi_{\Phi}(\cdot)$ $\sum_{j=1}^{q}{\phi_{\Phi_{j}}(\cdot)}:[0,+\infty)\to[0,+\infty)$ with $\phi_{\Phi}(r)\to 0$ as $r\to 0$, is a monotonic modulus of continuity of $\Phi$, such that ${\left\lVert\Phi[\mathbf{u}_{1}(\tau)](t)-\Phi[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}\leq\sum_{j=1}^{q}{\left\lVert\Phi_{j}[\mathbf{u}_{1}(\tau)](t)-\Phi_{j}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}\leq\phi_{\Phi}[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}}]$ for arbitrary $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in C_{0}([0,T];{\mathbb{R}^{p}})$.
Setting 1: Given two real continuous function spaces $C_{0}([0,T];{\mathbb{R}^{p}})$ and $C_{0}([0,T];{\mathbb{R}^{q}})$ with $L^{\infty}$-norm ${\left\lVert\cdot\right\rVert}_{L^{\infty}}$, a compact set $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$, and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$, for each $\mathbf{u}(t)\in K$, its corresponding image function $\mathbf{v}(t)\in C_{0}([0,T];{\mathbb{R}^{q}})$ is calculated by $\Phi$
@@
Informal Statement of Theorem 1: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})$ $\to C_{0}([0,T];{\mathbb{R}^{q}})$ satisfying Assumptions 1 to 6 with $\mathrm{\upphi}_{K}(t)=L_{K}[t,t,\dots,t]^{\top}$ and $L_{K}>0$, for every integer $M_{\it{\Gamma}}$ larger than an independent deterministic threshold and every positive integer $H_{\it{\Pi}}$, there exist two MLPs $\it{\Gamma}(\cdot)$ and $\it{\Pi}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for an arbitrary $\mathbf{u}(t)\in K$, the corresponding solution $\mathbf{y}(t)=[y_{1}(t),y_{2}(t),...,y_{q}(t)]^{\top}$ from Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)={\mathbf{x}}^{\prime}(0)=\mathbf{0}$ satisfies
@@
and ${\phi_{\Phi}}(\cdot)$ is the modulus of continuity of $\Phi$ in Assumption 6. $\it{\Gamma}(\cdot)$ has one hidden layer and the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}=O\left(pM_{{\it{\Gamma}}}\right)$, $w_{\it{\Gamma}}=O\left(pM_{{\it{\Gamma}}}\right)$, and $w_{{\it{\Gamma},{\mathrm{out}}}}=O\left(pM_{{\it{\Gamma}}}\right)$, respectively. For ${\it{\Pi}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Pi},{\mathrm{in}}}}=O\left(pM_{{\it{\Gamma}}}\right)$, $w_{\it{\Pi}}=O\left(qpM_{{\it{\Gamma}}}\right)$, and $w_{{\it{\Pi},{\mathrm{out}}}}=q$, respectively, and its depth is $h_{\it{\Pi}}=H_{\it{\Pi}}+1$.
@@
with ${\hat{\mathbf{x}}}(0)={{\hat{\mathbf{x}}}}^{\prime}(0)=\mathbf{0}$, is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on a domain $\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}\subset\mathbb{R}^{2r+p}$, where ${\hat{\mathbf{x}}}(t)=\left[{\hat{x}}_{1}(t),{\hat{x}}_{2}(t),...,{\hat{x}}_{r}(t)\right]^{\top}$, ${\hat{\mathbf{y}}}(t)=\left[{\hat{y}}_{1}(t),{\hat{y}}_{2}(t),...,{\hat{y}}_{q}(t)\right]^{\top}$, $B_{{\mathbf{x}}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}}$, $B_{{\mathbf{x}}^{\prime}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}^{\prime}}$, $\alpha$ is a positive coefficient, $B_{\hat{\mathbf{x}}}$ and $B_{\hat{\mathbf{x}}^{\prime}}$ are the bounds of ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$, respectively, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, and $B_{\beta{g}}$ is the stability bound of the system in Eq. (7). The functions $g(\cdot)=\left[g_{1}(\cdot),g_{2}(\cdot),...g_{r}(\cdot)\right]^{\top}$ and $h(\cdot)=\left[h_{1}(\cdot),h_{2}(\cdot),...h_{q}(\cdot)\right]^{\top}$ are Lipschitz continuous over ${\mathbb{R}}^{2r+p}$ and $\mathbb{R}^{r}$, respectively, with arbitrary finite-order derivatives. Then, for arbitrary positive errors $\varepsilon_{1}\leq\alpha/r$ and $\varepsilon_{2}$, there exist two one-hidden-layer MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for all $\mathbf{u}(t)\in K$, the corresponding solutions $\mathbf{y}(t)$ to Eq. (1) and ${\hat{\mathbf{y}}}(t)$ to Eq. (7) satisfy
@@ 3.1 Causal Continuous Operators
In this subsection, Lemma 1 derives an error bound for approximating all $\mathbf{u}(t)\in K$ using their initial values $\mathbf{u}(0)$ and a finite number of time-varying sine transform coefficients, combined with a set of corresponding deterministic base functions. Lemma 2 proves the existence of an MLP ${\it{\Gamma}}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, such that the solution $\mathbf{x}(t)$ to the first ODE in Eq. (1) corresponds to the sine transform coefficient vector of the input function $\mathbf{u}(t)$. Lemma 3 demonstrates that the target causal continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$ in Setting 1 can be approximated by a continuous function $\Psi\left[\mathbf{x}(t),\mathbf{u}(0),t\right]$, which depends on time $t$, $\mathbf{u}(0)$, and the solution $\mathbf{x}(t)$, and establishes an associated error bound. Lemma 4 determines the Lipschitz constants, linear moduli of continuity, and a compact argument domain of $\Psi\left[\mathbf{x}(t),\mathbf{u}(0),t\right]$. Finally, by approximating $\Psi\left[\mathbf{x}(t),\mathbf{u}(0),t\right]$ with an MLP ${\it{\Pi}}(\cdot)$ in Eq. (3) and combing its approximation error with the error bound from Lemma 3, Theorem 1 establishes an approximation bound for the neural oscillator in Eq. (1). The proofs for all lemmas and Theorem 1 in this subsection are presented in A.
Lemma 1: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ satisfying Assumptions 1 and 2, for an arbitrary $v$ satisfying $0<v<T$ and every positive integer M larger than one, there exist M frequencies $\omega_{1},\omega_{2},...,\omega_{M}\in\mathbb{R}$, M phases $\theta_{1},\theta_{2},...,\theta_{M}\in\mathbb{R}$, and M weights $\alpha_{1},\alpha_{2},...,\alpha_{M}\in\mathbb{R}$, such that
@@
${\varepsilon_{K}}=\left|{\mathbf{\upphi}_{K}(v)}\right|+\frac{{8T^{2}\left|{\mathbf{\upphi}_{K}(T)}\right|{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}M}},$ (11)
$c_{\rho,2}$ is an independent constant, and $\mathbf{\upphi}_{K}(t)$ is a modulus of continuity for all $\mathbf{u}(t)\in K$ in Assumption 2.
@@
${\varepsilon_{K}}=p{L_{K}}v+\frac{{8{T^{3}}p{L_{K}}{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}M}}.$ (12)
@@
${\varepsilon_{K}}=p{L_{K}}2T{\left({\frac{{{2}{c_{\rho,2}}}}{{{\pi^{2}}M}}}\right)^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}+\frac{{8{T^{3}}p{L_{K}}{c_{\rho,2}}}}{{{\pi^{2}}M}}\frac{1}{4T^{2}}{\left({\frac{{{\pi^{2}}M}}{{{2}{c_{\rho,2}}}}}\right)^{{2\mathord{\left/{\vphantom{23}}\right.\kern-1.2pt}3}}}=\frac{{54^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}p{L_{K}}Tc_{\rho,2}^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}}{{{\pi^{{2\mathord{\left/{\vphantom{23}}\right.\kern-1.2pt}3}}}{M^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}}},$ (13)
@@
$\frac{{{\rm{d}}{\varepsilon_{K}}}}{{{\rm{d}}v}}=p{L_{K}}-\frac{{16{T^{3}}p{L_{K}}{c_{\rho,2}}}}{{{\pi^{2}}{v^{3}}M}}=0\Rightarrow v=2T{\left({\frac{{{2}{c_{\rho,2}}}}{{{\pi^{2}}M}}}\right)^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}.$ (14)
Thus, for each positive integer M satisfying
$M>16{c_{\rho,2}}{\pi^{-2}},$ (15)
the approximation error $\varepsilon_{K}$ in Eq. (9) can be expressed in Eq. (13). The condition in Eq. (15) ensures that the underlying parameter $v$ is smaller than $T$.
Lemma 2: Given a compact subset K $\subset C_{0}([0,T];{\mathbb{R}^{p}})$ satisfying Assumptions 1 and 2, for every positive integer M larger than one and arbitrary M frequencies $\omega_{1}$, $\omega_{2}$,…, $\omega_{M}$ $\in\mathbb{R}$, there exists a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}=p(2M+1)$, $w_{\it{\Gamma}}=2pM$, and $w_{{\it{\Gamma},{\mathrm{out}}}}=pM$, respectively, such that for all $t\in[0,T]$ and all $\mathbf{u}(t)\in K$, the solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}$ from the second-order ODE in Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$ satisfies $\mathbf{x}(t)=[\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{1}),\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{2}),...,\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{M})]^{\top}$, where $\mathcal{L}_{t}\mathbf{u}(\omega)$ is defined in Eq. (10) in Lemma 1.
Lemma 3: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$ satisfying Assumptions 1, 2, 3, and 6, for an arbitrary positive $v<T$ and every positive integer M larger than one, there exists a continuous mapping $\Psi=[\psi_{1},\psi_{2},...,\psi_{q}]:{\mathbb{R}}^{p(M+1)}\times{[0,T]}\to{\mathbb{R}}^{q}$ and a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ specified in Lemma 2, such that for all $\mathbf{u}(t)\in K$, the solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}$ from the second-order ODE in Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$ satisfies
@@
Lemma 4: Under the conditions in Lemma 3 and Assumption 5, along with $\mathrm{\upphi}_{K}(t)=[L_{K}t,L_{K}t,...,L_{K}t]^{\top}$, $v$ satisfying Eq. (14), and an integer $M$ satisfying Eq. (15), the values of all arguments of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ in Lemma 3 are within a compact domain $\Omega_{M}\left(TB_{K},B_{K},T\right)=\left[-TB_{K},TB_{K}\right]^{pM}\times\left[-B_{K},B_{K}\right]^{p}\times\left[0,T\right]$ for all $\mathbf{u}(t)\in K$ and j=1,2,,qj=1,2,...,q, the Lipschtiz constant $L_{{\psi_{j}}}\left(M\right)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ with respect to its each argument is
$\begin{array}[]{l}{L_{{\psi_{j}}}\left(M\right)}=\max\left\{{{D_{{\Phi_{j}}}}+p{B_{K}}MT^{-1}c_{L}{{\phi^{\prime}}_{{\Phi_{j}}}}(0),\max\left[{\left({1+2M}\right),T^{-1}}\right]{{\phi^{\prime}}_{{\Phi_{j}}}}(0)}\right\},\end{array}$ (17)
and a linear modulus of continuity $\phi_{\psi_{j},M}(x)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ under $L^{2}$-norm is

### v2

@@ 2 Neural Oscillator
Assumption 6: Since ${{\Phi}}=[\Phi_{1},\Phi_{2}....,\Phi_{q}]$ is a causal and uniformly continuous operator mapping from $C_{0}([0,T];{\mathbb{R}}^{p})$ to $C_{0}([0,T];{\mathbb{R}}^{q})$, its $j^{\mathrm{th}}$ element $\Phi_{j}$ has a monotonic continuous modulus of continuity $\phi_{\Phi_{j}}(\cdot):[0,+\infty)\to[0,+\infty)$ with $\phi_{\Phi_{j}}(r)\to 0$ as $r\to 0$, such that ${\left\lVert\Phi_{j}[\mathbf{u}_{1}(\tau)](t)-\Phi_{j}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}$ $\leq\phi_{\Phi_{j}}[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}}]$ for arbitrary $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in C_{0}([0,T];{\mathbb{R}^{p}})$. $\phi_{\Phi}(\cdot)$ $\sum_{j=1}^{q}{\phi_{\Phi_{j}}(\cdot)}:[0,+\infty)\to[0,+\infty)$ with $\phi_{\Phi}(r)\to 0$ as $r\to 0$ is a monotonic modulus of continuity of $\Phi$, such that ${\left\lVert\Phi[\mathbf{u}_{1}(\tau)](t)-\Phi[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}\leq\sum_{j=1}^{q}{\left\lVert\Phi_{j}[\mathbf{u}_{1}(\tau)](t)-\Phi_{j}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}\leq\phi_{\Phi}[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}}]$ for arbitrary $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in C_{0}([0,T];{\mathbb{R}^{p}})$.
Setting 1: Given two real continuous function spaces $C_{0}([0,T];{\mathbb{R}^{p}})$ and $C_{0}([0,T];{\mathbb{R}^{q}})$ with the $L^{\infty}$-norm ${\left\lVert\cdot\right\rVert}_{L^{\infty}}$, a compact set $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$, and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$, for each $\mathbf{u}(t)\in K$, its corresponding image function $\mathbf{v}(t)\in C_{0}([0,T];{\mathbb{R}^{q}})$ is calculated by $\Phi$
@@
Informal Statement of Theorem 1: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})$ $\to C_{0}([0,T];{\mathbb{R}^{q}})$ satisfying Assumptions 1 to 6 with $\mathrm{\upphi}_{K}(t)=L_{K}[t,t,\dots,t]^{\top}$ and $L_{K}>0$, for every integer $M_{\it{\Gamma}}$ implicitly larger than an independent deterministic threshold and every positive integer $H_{\it{\Pi}}$, there exist two MLPs $\it{\Gamma}(\cdot)$ and $\it{\Pi}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for an arbitrary $\mathbf{u}(t)\in K$, the corresponding solution $\mathbf{y}(t)=[y_{1}(t),y_{2}(t),...,y_{q}(t)]^{\top}$ from Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)={\mathbf{x}}^{\prime}(0)=\mathbf{0}$ satisfies
@@
and ${\phi_{\Phi}}(\cdot)$ is the modulus of continuity of $\Phi$ in Assumption 6. $\it{\Gamma}(\cdot)$ has one hidden layer and the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}=2pM_{{\it{\Gamma}}}+p$, $w_{\it{\Gamma}}=2pM_{{\it{\Gamma}}}$, and $w_{{\it{\Gamma},{\mathrm{out}}}}=pM_{{\it{\Gamma}}}$, respectively. For ${\it{\Pi}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Pi},{\mathrm{in}}}}=pM_{{\it{\Gamma}}}+p+1$, $w_{\it{\Pi}}=q\left(pM_{{\it{\Gamma}}}+p+4\right)$, and $w_{{\it{\Pi},{\mathrm{out}}}}=q$, respectively, and its depth is $h_{\it{\Pi}}=H_{\it{\Pi}}+1$.
@@
with ${\hat{\mathbf{x}}}(0)={{\hat{\mathbf{x}}}}^{\prime}(0)=\mathbf{0}$, is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on a domain $\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}\subset\mathbb{R}^{2r+p}$, where ${\hat{\mathbf{x}}}(t)=\left[{\hat{x}}_{1}(t),{\hat{x}}_{2}(t),...,{\hat{x}}_{r}(t)\right]^{\top}$, ${\hat{\mathbf{y}}}(t)=\left[{\hat{y}}_{1}(t),{\hat{y}}_{2}(t),...,{\hat{y}}_{q}(t)\right]^{\top}$, $B_{{\mathbf{x}}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}}$, $B_{{\mathbf{x}}^{\prime}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}^{\prime}}$, $\alpha$ is a positive coefficient, $B_{\hat{\mathbf{x}}}$ and $B_{\hat{\mathbf{x}}^{\prime}}$ are the bounds of ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$, respectively, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, and $B_{\beta{g}}$ is the stability bound of the system in Eq. (7). $g(\cdot)=\left[g_{1}(\cdot),g_{2}(\cdot),\dots,g_{r}(\cdot)\right]^{\top}$ satisfies that for every compact domain $\Omega\subset$ ${\mathbb{R}}^{2r+p}$, there exists a Barron function extension $\tilde{g}_{\Omega}(\cdot)=g(\cdot)$ on $\Omega$ and the spectral Barron norms of all extensions are uniformly bounded. $h(\cdot)=\left[h_{1}(\cdot),h_{2}(\cdot),\dots,h_{q}(\cdot)\right]^{\top}$ satisfies that for every compact domain $\Omega\subset$ ${\mathbb{R}}^{r}$, there exists a Barron function extension $\tilde{h}_{\Omega}(\cdot)=h(\cdot)$ on $\Omega$ and the spectral Barron norms of all extensions are uniformly bounded. Then, for arbitrary positive errors $\varepsilon_{1}\leq\alpha/r$ and $\varepsilon_{2}$, there exist two one-hidden-layer MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for all $\mathbf{u}(t)\in K$, the corresponding solutions $\mathbf{y}(t)$ to Eq. (1) and ${\hat{\mathbf{y}}}(t)$ to Eq. (7) satisfy
@@
In the two theorems, the convergence norms are the $L^{\infty}$-norm. The activation functions $\sigma_{\it{\Gamma}}\left(\cdot\right)$ of ${\it{\Gamma}}(\cdot)$ and $\sigma_{\it{\Pi}}\left(\cdot\right)$ of ${\it{\Pi}}(\cdot)$ are $\sigma_{\mathrm{ReLU}}(\cdot)$. In the numerical studies in Section 4, $\sigma_{\mathrm{ReLU}}(\cdot)$ is adopted for ${\it{\Gamma}}(\cdot)$ and single-hidden-layer ${\it{\Pi}}(\cdot)$. For multi-hidden-layer ${\it{\Pi}}(\cdot)$, the Parametric Rectified Linear Unit (PReLU) $\sigma_{\text{PReLU}}(x)=\max(0,x)+\alpha\min(0,x)$, where $\alpha$ is a parameter learned during training (He et al., 2015), is employed to improve the training process.
@@ 3.1 Causal Continuous Operators
In this subsection, Lemma 1 derives an error bound for approximating all $\mathbf{u}(t)\in K$ using their initial values $\mathbf{u}(0)$ and a finite number of time-varying sine transform coefficients, combined with a set of corresponding deterministic base functions. Lemma 2 proves the existence of an MLP ${\it{\Gamma}}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, such that the solution $\mathbf{x}(t)$ of the neural ODE in Eq. (1) corresponds to the sine transform coefficient vector of the input function $\mathbf{u}(t)$. Lemma 3 demonstrates that the target causal continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$ in Setting 1 can be approximated by a continuous function $\Psi\left[\mathbf{x}(t),\mathbf{u}(0),t\right]$, which depends on time $t$, $\mathbf{u}(0)$, and the solution $\mathbf{x}(t)$, and establishes an associated error bound. Lemma 4 determines the Lipschitz constants, linear moduli of continuity, and a compact argument domain of $\Psi\left[\mathbf{x}(t),\mathbf{u}(0),t\right]$. Finally, by approximating $\Psi\left[\mathbf{x}(t),\mathbf{u}(0),t\right]$ with an MLP ${\it{\Pi}}(\cdot)$ in Eq. (3) and combing its approximation error with the error bound from Lemma 3, Theorem 1 establishes an approximation bound for the neural oscillator in Eq. (1). The proofs for all lemmas and Theorem 1 in this subsection are presented in A.
Lemma 1: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ satisfying Assumptions 1 and 2, for an arbitrary $v$ satisfying $0<v<T$ and every positive integer M larger than two, there exist M frequencies $\omega_{1},\omega_{2},...,\omega_{M}\in\mathbb{R}$, M phases $\theta_{1},\theta_{2},...,\theta_{M}\in\mathbb{R}$, and M weights $\alpha_{1},\alpha_{2},...,\alpha_{M}\in\mathbb{R}$, such that
@@
${\varepsilon_{K}}=\left|{\mathbf{\upphi}_{K}(v)}\right|+\frac{c_{K}^{r+1}T^{r}\left|{\mathbf{\upphi}_{K}(T)}\right|\left(r!\right)^{2}\ln{M}}{v^{r}M^{r}},$ (11)
$c_{K}>1$ is an independent constant, $r$ can be an arbitrary positive integer, and $\mathbf{\upphi}_{K}(t)$ is a modulus of continuity for all $\mathbf{u}(t)\in K$ in Assumption 2.
@@
${\varepsilon_{K}}=p{L_{K}}v+\frac{c_{K}^{r+1}pL_{k}T^{r+1}\left(r!\right)^{2}\ln{M}}{v^{r}M^{r}}.$ (12)
@@
${\varepsilon_{K}}=p{L_{K}}\left[r^{1/(r+1)}+r^{-r/(r+1)}\right]\frac{c_{K}T\left(r!\right)^{2/(r+1)}\left(\ln{M}\right)^{1/(r+1)}}{M^{r/(r+1)}}$ (13)
[…]
@@
$\frac{{{\rm{d}}{\varepsilon_{K}}}}{{{\rm{d}}v}}=p{L_{K}}-\frac{rc_{K}^{r+1}pL_{k}T^{r+1}\left(r!\right)^{2}\ln{M}}{v^{r+1}M^{r}}=0\Rightarrow v=\frac{r^{1/(r+1)}c_{K}T\left(r!\right)^{2/(r+1)}\left(\ln{M}\right)^{1/(r+1)}}{M^{r/(r+1)}},$ (14)
where in Step (a) $r=\lfloor\ln{M}\rfloor$ and $M^{1/(1+r)}\leq M^{1/\lfloor\ln{M}\rfloor}=e^{\ln{M}/\lfloor\ln{M}\rfloor}<e^{2}$ for $M>2$ are utilized. The appropriate ranges of $v$ and $M$ are
$\left\{\begin{aligned} &\frac{T}{M}\leq v\leq c_{K}T\frac{r^{(2r+1)/(r+1)}\left(\ln{M}\right)^{1/(r+1)}}{M^{r/\left(1+r\right)}}\\ &c_{K}\leq\frac{M^{r/\left(1+r\right)}}{r^{(2r+1)/(r+1)}\left(\ln{M}\right)^{1/(r+1)}}\end{aligned}.\right$. (15)
where $r=\lfloor\ln{M}\rfloor$. Thus, for each integer M satisfying Eq. (15), the approximation error $\varepsilon_{K}$ in Eq. (9) can be expressed by the result in Eq. (13). The condition of $M$ in Eq. (15) ensures that the underlying parameter $v$ is smaller than $T$.
Lemma 2: Given a compact subset K $\subset C_{0}([0,T];{\mathbb{R}^{p}})$ satisfying Assumptions 1 and 2, for every positive integer M larger than two and arbitrary M frequencies $\omega_{1}$, $\omega_{2}$,…, $\omega_{M}$ $\in\mathbb{R}$, there exists a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}=p(2M+1)$, $w_{\it{\Gamma}}=2pM$, and $w_{{\it{\Gamma},{\mathrm{out}}}}=pM$, respectively, such that for all $t\in[0,T]$ and all $\mathbf{u}(t)\in K$, the solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}$ from the second-order ODE in Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$ satisfies $\mathbf{x}(t)=[\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{1}),\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{2}),...,\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{M})]^{\top}$, where $\mathcal{L}_{t}\mathbf{u}(\omega)$ is defined in Eq. (10) in Lemma 1.
Lemma 3: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$ satisfying Assumptions 1, 2, 3, and 6, for an arbitrary positive $v<T$ and every positive integer M larger than two, there exists a continuous mapping $\Psi=[\psi_{1},\psi_{2},...,\psi_{q}]:{\mathbb{R}}^{p(M+1)}\times{[0,T]}\to{\mathbb{R}}^{q}$ and a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ specified in Lemma 2, such that for all $\mathbf{u}(t)\in K$, the solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}$ from the second-order ODE in Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$ satisfies
@@
Lemma 4: Under the conditions in Lemma 3 and Assumption 5, along with $\mathrm{\upphi}_{K}(t)=[L_{K}t,L_{K}t,...,L_{K}t]^{\top}$, $v$ satisfying Eq. (14), and an integer $M$ satisfying Eq. (15), the values of all arguments of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ in Lemma 3 are within a compact domain $\Omega_{M}\left(TB_{K},B_{K},T\right)=\left[-TB_{K},TB_{K}\right]^{pM}\times\left[-B_{K},B_{K}\right]^{p}\times\left[0,T\right]$ for all $\mathbf{u}(t)\in K$ and j=1,2,,qj=1,2,...,q, the Lipschitz constant $L_{{\psi_{j}}}\left(M\right)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ with respect to its each argument is
${L_{{\psi_{j}}}\left(M\right)}=\max\left\{{{D_{{\Phi_{j}}}}+p{B_{K}}M^{3}T^{-1}c_{L}{{\phi^{\prime}}_{{\Phi_{j}}}}(0),\max\left[{\left({1+2M}\right),T^{-1}}\right]{{\phi^{\prime}}_{{\Phi_{j}}}}(0)}\right\},$ (17)
and a linear modulus of continuity $\phi_{\psi_{j},M}(x)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ under the $L^{2}$-norm is